OpenStudy (anonymous):

in triangle abc the measure of angle b is 36 more than three times the measure of angle a. the measure of angle c is 54 more than the measure of angle a. find the measure of each angle

OpenStudy (anonymous):

b = 36 + 3a c = 54 + a a + b + c = 180 a + 36 + 3a + 54 + a = 180 5a = 180 - 90 5a = 90 a = 16 b = 36 + 3(16) = 36 + 48 = 84 c = 54 + 16 = 70

OpenStudy (anonymous):

a + (3 a + 36) + (a + 54) = 180 a = 18, b = 90, c = 72

OpenStudy (anonymous):

yeah robtobey is right,....i make mistake in my work,...90 / 5 is 18 for the measure of a... sorry,...applause for u rob...